# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v3

@@ 8 Paraphrased sentence:
gemma-2b baseline zero-shot 0.563 74.08 21.41 4.51 29.17 32.14 67.69 44.19
1-shot 0.558 74.54 21.34 4.13 32.37 31.49 63.27 45.16
5-shot 0.474 81.61 15.89 2.51 26.28 30.19 55.78 48.06
10-shot 0.421 82.49 16.01 1.51 23.72 23.38 48.64 48.06
sft zero-shot 0.268 84.91 14.20 0.89 13.14 18.51 26.87 35.48
gemma-1.1- 2b-it baseline zero-shot 0.612 59.20 32.25 8.55 31.09 28.25 55.44 35.48
1-shot 0.565 66.76 28.47 4.77 32.37 29.55 55.10 39.35
5-shot 0.471 66.50 27.44 6.06 25.64 24.35 47.96 31.94
10-shot 0.468 68.98 26.44 4.58 26.28 22.40 49.66 35.48
sft zero-shot 0.250 81.27 17.78 0.95 11.86 15.26 26.87 30.00
gemma-7b baseline zero-shot 0.546 72.28 24.24 3.48 29.49 32.14 59.18 42.90
1-shot 0.535 77.74 18.10 4.15 30.77 31.82 57.82 51.61
5-shot 0.437 84.39 13.79 1.81 25.64 26.95 47.96 51.94
10-shot 0.369 84.95 13.33 1.72 22.44 23.05 35.71 48.06
sft zero-shot 0.203 81.64 16.80 1.56 10.26 11.36 18.03 28.71
gemma-1.1- 7b-it baseline zero-shot 0.337 55.06 38.35 6.59 9.29 14.29 26.53 26.77
1-shot 0.310 69.57 26.09 4.35 14.10 13.31 26.53 35.16
5-shot 0.279 73.01 25.00 1.99 13.14 13.64 20.07 37.10
10-shot 0.232 73.72 24.91 1.37 10.58 7.79 15.31 36.77
sft zero-shot 0.214 80.74 17.78 1.48 10.90 11.04 17.35 31.94
Table 15: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
@@
gemma-2b baseline zero-shot 0.803 274 33.29 33.42 33.29 20.51 23.05 21.43 15.48
1-shot 0.750 0 31.18 29.6 39.22 22.44 26.62 23.47 23.87
5-shot 0.722 0 30.63 23.82 45.55 20.19 23.70 28.23 19.35
10-shot 0.736 0 29.42 22.74 47.84 21.47 23.70 23.81 20.32
sft zero-shot 0.741 6 32.54 34.48 32.97 22.44 27.92 25.51 22.90
gemma-1.1- 2b-it baseline zero-shot 0.738 4 34.23 27.86 37.90 24.04 28.25 25.85 25.48
1-shot 0.654 0 46.67 18.91 34.42 29.81 35.71 30.61 29.68
5-shot 0.631 0 48.68 14.59 36.73 32.05 37.34 27.55 29.35
10-shot 0.660 0 43.87 11.06 45.07 29.49 34.42 25.85 29.35
sft zero-shot 0.994 1,239 85.71 0 14.29 0.96 0.97 1.02 0.97
gemma-7b baseline zero-shot 0.612 195 57.89 9.88 32.24 17.31 19.81 44.90 28.06
1-shot 0.351 0 69.07 12.19 18.74 20.51 14.61 40.82 24.84
5-shot 0.307 0 77.52 15.50 6.98 22.76 17.86 35.03 22.90
10-shot 0.288 0 81.54 13.50 4.96 24.36 16.23 31.29 25.16
sft zero-shot 0.278 1 70.57 12.57 16.86 11.54 14.29 28.91 26.45
gemma-1.1- 7b-it baseline zero-shot 0.355 0 64.51 20.09 15.40 16.99 15.26 40.14 22.90
1-shot 0.307 0 63.05 15.76 21.19 16.03 11.36 34.01 19.03
5-shot 0.283 0 63.87 18.77 17.37 16.99 9.42 29.59 19.03
10-shot 0.289 0 65.93 17.31 16.76 19.23 8.77 30.61 20.32
sft zero-shot 0.269 2 76.26 16.91 6.820 22.12 14.61 27.21 20.32
Table 16: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
@@
Llama-3.2- 3B baseline zero-shot 0.556 70.04 23.11 6.85 24.68 30.19 62.59 44.19
1-shot 0.519 75.99 20.49 3.52 25.32 30.19 60.54 47.42
5-shot 0.419 78.41 18.94 2.65 20.19 25.32 46.60 43.87
10-shot 0.401 80.20 16.83 2.97 22.44 24.03 43.20 43.23
sft zero-shot 0.267 82.79 15.43 1.78 14.42 17.53 25.85 33.55
Llama-3.2- 3B-Instruct baseline zero-shot 0.528 72.67 23.87 3.45 23.72 29.55 50.68 54.84
1-shot 0.502 75.36 23.22 1.42 28.85 27.27 48.98 51.29
5-shot 0.441 76.08 22.84 1.08 25.64 26.62 39.80 46.45
10-shot 0.410 76.79 21.86 1.35 25.64 24.68 35.03 44.52
sft zero-shot 0.255 83.80 14.95 1.25 12.82 13.31 27.55 34.52
Llama-3.1- 8B baseline zero-shot 0.562 70.62 21.33 8.05 25.64 30.84 64.29 43.87
1-shot 0.507 77.78 18.15 4.07 25.96 29.87 59.18 48.39
5-shot 0.393 83.87 14.31 1.81 20.83 22.40 47.62 45.81
10-shot 0.358 84.51 13.72 1.77 20.19 23.05 40.82 41.29
sft zero-shot 0.203 85.55 12.50 1.95 8.97 11.69 18.71 32.26
Llama-3.1- 8B-Instruct baseline zero-shot 0.488 70.57 26.18 3.25 24.36 25.97 54.76 37.74
1-shot 0.419 73.86 24.43 1.70 22.76 25.97 44.56 34.84
5-shot 0.347 80.55 18.76 0.69 18.91 19.16 39.80 37.74
10-shot 0.329 81.20 18.31 0.48 19.55 20.45 34.01 36.45
sft zero-shot 0.244 86.04 12.34 1.62 12.18 12.66 28.23 33.87
Table 17: Error rates, incorrect choice distributions, and local negation confusion rates for the LLaMA family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
@@
Llama-3.2- 3B baseline zero-shot 0.688 0 42.21 25.95 31.83 22.12 33.12 32.65 31.94
1-shot 0.570 0 62.59 14.74 22.67 31.09 32.14 45.58 38.71
5-shot 0.548 0 74.10 14.33 11.58 34.62 35.06 50.34 47.74
10-shot 0.500 0 78.13 13.00 8.87 35.90 31.82 43.88 49.68
sft zero-shot 0.658 0 39.40 33.01 27.59 23.40 31.17 27.89 24.52
Llama-3.2- 3B-Instruct baseline zero-shot 0.474 8 63.90 10.34 25.76 19.55 28.25 60.20 16.77
1-shot 0.461 0 71.60 12.56 15.83 23.40 25.97 61.22 26.77
5-shot 0.472 0 74.62 10.92 14.45 29.81 28.25 63.27 25.16
10-shot 0.467 0 73.85 10.36 15.79 33.01 25.00 62.24 23.23
sft zero-shot 0.437 87 61.85 7.11 31.03 11.22 18.51 43.54 21.61
Llama-3.1- 8B baseline zero-shot 0.541 0 47.95 6.74 45.31 23.08 24.03 41.50 19.03
1-shot 0.454 0 67.66 4.55 27.80 27.56 28.25 47.62 23.87
5-shot 0.362 0 78.95 6.58 14.47 30.45 21.75 40.82 25.16
10-shot 0.316 0 82.46 6.02 11.53 30.77 21.75 31.97 23.23
sft zero-shot 0.444 48 58.79 27.34 13.87 20.83 20.13 37.41 20.65
Llama-3.1-8B- Instruct baseline zero-shot 0.256 12 76.53 15.11 8.36 16.67 18.83 19.05 23.23
1-shot 0.290 0 71.51 21.37 7.12 19.23 18.18 29.25 19.03
5-shot 0.269 0 74.04 21.83 4.13 20.83 16.88 25.17 19.35
10-shot 0.290 0 80.00 14.52 5.48 25.00 20.78 26.87 22.90
sft zero-shot 0.254 31 72.32 19.03 8.65 15.06 17.86 21.09 14.52
Table 18: Error rates, incorrect choice distributions, and local negation confusion rates for the LLaMA family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
@@
Mistral-7B- v0.3 baseline zero-shot 0.575 76.97 18.48 4.55 26.60 32.14 66.33 58.39
1-shot 0.532 78.39 17.59 4.02 26.92 33.12 59.52 53.23
5-shot 0.439 83.57 14.98 1.44 22.12 25.65 51.70 52.58
10-shot 0.401 84.39 13.83 1.78 21.15 24.68 47.28 47.10
sft zero-shot 0.227 88.81 9.09 2.10 8.65 13.31 29.93 31.61
Mistral-7B- Instruct-v0.3 baseline zero-shot 0.381 67.29 30.83 1.88 13.46 18.51 30.95 42.90
1-shot 0.365 72.61 26.09 1.30 15.38 20.45 34.69 39.03
5-shot 0.331 72.66 26.86 0.48 14.10 15.58 31.97 37.74
10-shot 0.304 73.63 25.85 0.52 12.18 14.29 30.61 35.48
sft zero-shot 0.236 85.52 13.80 0.67 10.26 12.01 27.21 33.87
Table 19: Error rates, incorrect choice distributions, and local negation confusion rates for the Mistral family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
@@
Mistral-7B- v0.3 baseline zero-shot 0.624 351 53.21 21.56 25.23 16.99 18.18 26.53 14.52
1-shot 0.410 0 67.12 16.63 16.25 30.77 24.68 28.91 29.03
5-shot 0.306 0 83.16 11.92 4.92 26.28 24.35 24.49 29.68
10-shot 0.272 0 85.42 8.45 6.12 23.40 22.73 21.09 28.39
sft zero-shot 0.614 0 40.05 29.46 30.49 22.44 27.92 28.57 22.58
Mistral-7B- Instruct-v0.3 baseline zero-shot 0.358 1 65.56 25.78 8.67 27.24 23.05 27.55 18.71
1-shot 0.370 0 61.03 34.90 4.07 27.24 20.78 26.19 19.03
5-shot 0.335 0 68.96 29.86 1.18 28.85 19.81 26.19 20.32
10-shot 0.331 0 71.22 27.10 1.68 28.53 22.08 25.51 20.97
sft zero-shot 0.351 1 67.57 21.32 11.11 23.08 23.05 33.67 18.06
Table 20: Error rates, incorrect choice distributions, and local negation confusion rates for the Mistral family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)

### v4

@@ 8 Paraphrased sentence:
gemma-2-2b baseline zero-shot 0.564 74.96 19.41 5.63 25.64 32.47 63.27 53.87
1-shot 0.523 76.52 19.55 3.94 28.85 30.84 58.84 47.42
5-shot 0.420 81.47 17.01 1.51 21.47 25.00 51.02 44.19
10-shot 0.399 81.71 16.30 1.99 21.47 21.43 46.94 45.16
after SFT zero-shot 0.219 82.25 17.03 0.72 11.86 12.99 26.19 23.55
gemma-2-2b-it baseline zero-shot 0.485 75.29 22.09 2.62 22.12 23.38 48.64 56.77
1-shot 0.480 77.36 21.32 1.32 29.17 24.68 50.00 49.68
5-shot 0.427 78.07 20.07 1.86 23.08 23.05 46.94 44.84
10-shot 0.408 77.24 21.60 1.17 21.47 20.78 41.50 46.45
after SFT zero-shot 0.259 82.87 15.60 1.53 11.86 15.58 28.57 32.90
gemma-2-9b baseline zero-shot 0.530 75.30 19.61 5.09 24.36 31.49 61.56 48.06
1-shot 0.500 78.73 17.62 3.65 26.28 29.87 58.84 48.06
5-shot 0.431 83.79 14.00 2.21 22.44 27.27 50.00 49.68
10-shot 0.402 84.22 14.60 1.18 21.15 27.27 44.22 47.42
after SFT zero-shot 0.229 85.47 12.11 2.42 10.26 13.96 22.45 34.19
gemma-2-9b-it baseline zero-shot 0.493 76.01 22.06 1.93 24.68 27.92 52.72 49.68
1-shot 0.439 77.26 21.66 1.08 24.36 24.68 46.94 44.52
5-shot 0.352 78.38 20.72 0.90 17.31 17.86 31.97 46.77
10-shot 0.302 77.95 21.00 1.05 14.10 16.23 26.19 40.65
after SFT zero-shot 0.233 87.41 10.88 1.70 9.29 14.94 26.19 33.87
Table 20: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma2 family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
@@
gemma-2-2b baseline zero-shot 0.765 33.20 26.66 40.15 22.12 29.55 29.25 23.87
1-shot 0.668 40.38 13.54 46.08 21.79 26.62 34.01 29.03
5-shot 0.623 53.69 12.47 33.84 27.88 34.42 40.48 35.48
10-shot 0.599 58.81 10.86 30.33 31.09 35.39 43.20 35.81
after SFT zero-shot 0.740 32.48 34.62 32.90 22.76 27.92 25.51 22.90
gemma-2-2b-it baseline zero-shot 0.496 74.88 8.80 16.32 37.18 26.30 66.67 24.19
1-shot 0.496 70.93 11.34 17.73 30.45 30.84 55.78 29.03
5-shot 0.473 67.67 12.56 19.77 31.09 33.12 43.88 24.52
10-shot 0.467 69.10 12.39 18.51 34.62 31.82 42.18 24.84
after SFT zero-shot 0.457 68.75 6.77 24.48 29.17 30.84 50.68 19.68
gemma-2-9b baseline zero-shot 0.532 62.89 7.60 29.51 36.22 29.87 57.82 15.16
1-shot 0.472 62.35 4.71 32.94 34.94 24.35 48.30 14.52
5-shot 0.402 71.79 6.11 22.09 33.01 23.70 45.92 17.10
10-shot 0.378 71.43 4.62 23.95 33.33 20.78 40.82 16.77
after SFT zero-shot 0.249 64.01 16.24 19.75 21.15 11.36 23.81 9.68
gemma-2-9b-it baseline zero-shot 0.447 73.00 22.02 4.97 39.42 22.40 51.70 21.61
1-shot 0.370 80.69 13.95 5.36 35.26 22.73 42.86 22.58
5-shot 0.286 83.66 11.63 4.71 28.85 17.53 31.63 20.97
10-shot 0.272 82.51 12.54 4.96 28.85 17.53 28.23 18.06
after SFT zero-shot 0.186 73.62 21.70 4.68 11.86 10.71 25.51 9.03
Table 21: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma2 family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
@@
gemma-2-2b baseline zero-shot 0.562 74.89 19.04 6.06 26.60 34.74 60.88 52.26
1-shot 0.507 76.37 19.41 4.23 26.60 30.84 56.46 46.45
5-shot 0.425 82.09 16.04 1.87 20.83 25.97 52.72 45.16
10-shot 0.392 81.98 15.99 2.02 20.19 22.40 46.26 44.19
after SFT zero-shot 0.232 82.59 16.72 0.68 11.22 13.64 27.21 27.42
gemma-2-2b-it baseline zero-shot 0.495 78.37 18.75 2.88 24.68 27.60 54.76 53.55
1-shot 0.493 79.74 18.81 1.45 29.17 28.25 52.38 52.90
5-shot 0.450 80.81 17.61 1.58 25.64 24.35 51.70 49.03
10-shot 0.422 80.26 18.42 1.32 22.76 22.40 46.60 48.39
after SFT zero-shot 0.267 81.90 16.62 1.48 12.50 14.61 32.65 30.97
gemma-2-9b baseline zero-shot 0.531 74.63 20.60 4.78 25.00 33.77 62.59 43.23
1-shot 0.487 78.50 17.59 3.91 25.96 30.52 59.18 42.90
5-shot 0.425 83.21 14.37 2.43 21.47 25.65 51.36 48.06
10-shot 0.393 83.23 15.35 1.41 20.83 25.97 43.54 44.84
after SFT zero-shot 0.228 86.06 11.85 2.09 8.97 14.61 23.81 33.55
gemma-2-9b-it baseline zero-shot 0.458 76.95 21.14 1.91 18.59 27.60 56.80 43.23
1-shot 0.401 77.62 21.58 0.79 20.51 20.78 45.24 42.26
5-shot 0.345 78.85 20.23 0.92 16.35 17.53 33.67 44.84
10-shot 0.308 78.87 19.85 1.29 14.10 17.21 26.53 42.26
after SFT zero-shot 0.236 85.91 11.41 2.68 10.26 13.64 26.87 33.23
Table 22: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma2 family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
@@
gemma-2-2b baseline zero-shot 0.724 35.93 29.46 34.61 23.72 31.49 29.25 22.90
1-shot 0.660 42.91 15.02 42.07 22.76 28.90 35.37 30.00
5-shot 0.629 53.47 12.48 34.05 28.21 34.74 40.82 35.16
10-shot 0.612 57.64 11.53 30.83 32.05 34.42 41.84 37.42
after SFT zero-shot 0.741 32.55 34.58 32.87 22.76 27.92 25.51 23.23
gemma-2-2b-it baseline zero-shot 0.531 82.06 7.47 10.46 42.95 38.96 67.69 30.97
1-shot 0.539 70.84 8.84 20.32 32.37 31.82 63.27 30.97
5-shot 0.508 64.06 12.81 23.12 30.13 32.47 45.92 26.13
10-shot 0.504 61.64 12.11 26.26 31.41 29.22 43.20 24.84
after SFT zero-shot 0.337 72.00 10.12 17.88 21.15 25.65 35.71 18.06
gemma-2-9b baseline zero-shot 0.534 73.70 8.47 17.83 42.63 32.47 66.33 21.94
1-shot 0.476 64.83 6.50 28.67 36.22 25.32 50.00 16.45
5-shot 0.422 70.11 5.45 24.44 33.65 22.73 48.30 18.06
10-shot 0.403 69.88 3.94 26.18 34.62 20.45 44.22 17.42
after SFT zero-shot 0.203 75.00 11.72 13.28 22.12 12.01 16.33 12.26
gemma-2-9b-it baseline zero-shot 0.431 83.27 11.58 5.15 35.90 23.70 63.61 26.13
1-shot 0.361 83.30 10.77 5.93 35.58 21.75 43.54 23.55
5-shot 0.305 84.11 10.42 5.47 30.45 21.43 32.31 21.61
10-shot 0.284 82.12 12.85 5.03 30.13 18.51 28.57 19.03
after SFT zero-shot 0.195 69.92 17.07 13.01 16.67 13.96 12.59 12.90
Table 23: Error rates, incorrect choice distributions, and local negation confusion rates for the Gemma2 family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
@@
Qwen3-0.6B-Base baseline zero-shot 0.547 72.32 23.48 4.20 29.49 34.74 50.68 48.71
1-shot 0.487 74.27 22.15 3.58 31.09 28.90 46.60 42.90
5-shot 0.412 77.07 19.46 3.47 23.40 21.75 43.54 42.58
10-shot 0.390 79.27 18.90 1.83 22.76 22.08 40.82 42.26
after SFT zero-shot 0.292 83.97 15.22 0.82 18.59 17.53 30.27 34.84
Qwen3-0.6B baseline zero-shot 0.584 74.46 20.52 5.03 26.92 32.14 64.63 56.45
1-shot 0.571 76.81 19.17 4.03 29.81 34.42 62.24 55.16
5-shot 0.515 79.82 16.95 3.24 25.96 28.57 57.48 58.06
10-shot 0.470 80.74 15.71 3.55 25.00 25.97 48.64 57.10
after SFT zero-shot 0.355 84.82 13.84 1.34 20.51 20.78 38.10 45.16
Qwen3-1.7B-Base baseline zero-shot 0.519 73.70 22.02 4.28 28.21 34.74 56.80 38.71
1-shot 0.467 74.87 22.92 2.21 27.88 30.19 50.34 36.45
5-shot 0.416 76.57 21.14 2.29 22.12 26.95 42.86 40.00
10-shot 0.396 79.56 18.84 1.60 23.72 24.35 41.84 40.32
after SFT zero-shot 0.301 89.47 9.74 0.79 19.23 24.35 36.05 31.94
Qwen3-1.7B baseline zero-shot 0.602 67.46 24.64 7.91 28.53 37.99 65.31 36.77
1-shot 0.601 71.50 24.01 4.49 31.41 34.74 64.29 47.74
5-shot 0.517 76.61 20.64 2.75 26.92 28.25 58.84 50.65
10-shot 0.483 78.33 19.70 1.97 27.88 24.35 54.08 50.32
after SFT zero-shot 0.332 83.77 14.56 1.67 15.71 24.68 37.41 37.42
Qwen3-4B-Base baseline zero-shot 0.546 76.34 20.46 3.19 31.73 32.47 62.59 46.13
1-shot 0.515 74.73 22.03 3.24 30.13 30.52 54.42 44.19
5-shot 0.467 74.19 22.41 3.40 26.28 27.27 48.64 41.29
10-shot 0.443 76.21 21.65 2.15 25.64 25.65 47.96 40.65
after SFT zero-shot 0.267 89.61 10.09 0.30 16.03 20.13 30.95 31.94
Qwen3-4B baseline zero-shot 0.590 73.66 20.83 5.51 31.73 40.58 68.37 39.68
1-shot 0.534 76.97 18.13 4.90 35.26 33.77 62.24 39.03
5-shot 0.435 80.66 17.34 2.01 25.00 25.97 52.72 41.61
10-shot 0.408 81.91 16.34 1.75 23.72 26.95 48.30 39.35
after SFT zero-shot 0.285 85.79 12.81 1.39 16.03 18.83 24.83 40.97
Qwen3-8B-Base baseline zero-shot 0.526 73.76 22.47 3.77 27.56 33.12 61.56 38.71
1-shot 0.502 72.67 23.85 3.48 27.24 28.57 56.12 39.35
5-shot 0.420 73.91 22.68 3.40 20.83 23.05 44.22 40.32
10-shot 0.401 77.62 20.40 1.98 22.44 25.32 40.14 40.65
after SFT zero-shot 0.297 85.07 13.60 1.33 16.67 19.81 35.03 33.23
Qwen3-8B baseline zero-shot 0.558 71.59 22.87 5.54 27.24 34.74 65.31 38.71
1-shot 0.501 77.22 18.20 4.59 32.37 32.14 61.22 34.84
5-shot 0.436 80.73 16.73 2.55 26.92 30.84 48.98 39.03
10-shot 0.397 81.20 17.00 1.80 25.32 26.95 44.56 36.45
after SFT zero-shot 0.278 83.48 15.10 1.42 14.74 16.56 31.97 32.90
Qwen3-14B-Base baseline zero-shot 0.512 70.08 23.57 6.36 24.68 30.84 53.74 39.35
1-shot 0.450 72.01 23.06 4.93 24.68 26.62 45.58 37.42
5-shot 0.380 75.57 21.50 2.92 20.83 22.73 37.76 37.42
10-shot 0.347 76.89 21.05 2.06 20.51 20.13 31.97 37.42
Qwen3-14B baseline zero-shot 0.560 72.52 21.10 6.37 27.88 32.47 65.31 42.90
1-shot 0.500 78.92 16.64 4.44 32.37 31.17 59.18 40.97
5-shot 0.429 80.41 17.19 2.40 25.64 26.95 47.28 42.90
10-shot 0.381 82.29 15.62 2.08 23.72 25.00 36.05 44.52
Qwen3-32B baseline zero-shot 0.531 76.68 17.64 5.68 27.24 34.42 67.01 40.32
1-shot 0.462 77.49 17.70 4.81 26.28 28.25 55.78 38.06
5-shot 0.347 81.46 15.79 2.75 18.91 22.73 39.80 35.48
10-shot 0.316 83.17 14.32 2.51 19.87 20.78 32.31 35.48
Table 24: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen3 family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
@@
Qwen3-0.6B-Base baseline zero-shot 0.651 50.55 7.06 42.39 25.32 37.99 39.80 32.90
1-shot 0.603 55.26 6.84 37.89 25.64 37.01 48.98 26.45
5-shot 0.577 55.57 3.30 41.13 27.56 33.12 47.28 24.84
10-shot 0.525 60.12 3.78 36.10 29.81 30.52 45.92 24.52
after SFT zero-shot 0.521 67.43 11.42 21.16 32.37 39.94 42.18 30.65
Qwen3-0.6B baseline zero-shot 0.562 69.96 7.90 22.14 37.82 38.64 54.08 32.26
1-shot 0.532 59.31 14.75 25.93 26.28 29.87 47.28 27.42
5-shot 0.493 71.66 5.80 22.54 32.37 30.52 54.76 28.71
10-shot 0.469 69.88 6.60 23.52 32.37 29.87 49.32 24.19
after SFT zero-shot 0.445 67.02 10.34 22.64 24.68 27.27 47.96 23.87
Qwen3-1.7B-Base baseline zero-shot 0.474 65.22 9.87 24.92 31.41 23.70 52.04 21.29
1-shot 0.444 72.50 13.57 13.93 33.97 30.84 38.78 29.35
5-shot 0.361 75.82 12.53 11.65 29.81 25.32 36.39 21.61
10-shot 0.313 76.14 11.42 12.44 25.96 20.78 34.69 17.10
after SFT zero-shot 0.507 58.84 12.83 28.33 33.97 26.30 43.54 19.68
Qwen3-1.7B baseline zero-shot 0.468 70.85 15.76 13.39 26.92 31.82 46.26 32.26
1-shot 0.451 67.31 14.24 18.45 24.04 28.57 47.62 25.81
5-shot 0.424 72.47 11.80 15.73 26.60 23.70 51.70 25.48
10-shot 0.421 72.13 9.98 17.89 25.96 23.70 53.40 23.23
after SFT zero-shot 0.392 68.83 22.27 8.91 30.13 27.92 30.27 22.90
Qwen3-4B-Base baseline zero-shot 0.398 78.69 15.74 5.58 38.14 22.40 45.58 23.55
1-shot 0.386 77.82 19.30 2.87 40.71 29.22 36.73 17.42
5-shot 0.350 75.74 19.27 4.99 40.38 26.95 24.15 17.42
10-shot 0.336 77.36 17.45 5.19 39.10 26.95 24.49 16.45
after SFT zero-shot 0.291 89.37 8.45 2.18 32.05 18.18 43.20 14.52
Qwen3-4B baseline zero-shot 0.439 82.28 15.55 2.17 38.46 30.19 56.46 24.52
1-shot 0.375 80.97 17.34 1.69 33.97 25.32 46.26 20.32
5-shot 0.321 84.20 12.35 3.46 32.69 23.70 33.67 21.61
10-shot 0.301 83.64 13.19 3.17 30.77 24.68 26.87 21.29
after SFT zero-shot 0.278 87.14 9.43 3.43 28.53 22.73 37.76 11.29
Qwen3-8B-Base baseline zero-shot 0.310 79.03 13.81 7.16 29.17 20.78 33.33 18.06
1-shot 0.313 76.96 13.42 9.62 31.73 26.95 23.81 16.77
5-shot 0.254 76.25 15.00 8.75 28.21 21.10 16.67 13.55
10-shot 0.250 78.41 13.02 8.57 27.56 20.45 16.67 15.81
after SFT zero-shot 0.232 72.7 20.82 6.48 19.55 17.21 18.71 14.19
Qwen3-8B baseline zero-shot 0.411 84.17 12.55 3.28 41.67 31.82 49.32 20.32
1-shot 0.326 82.00 13.14 4.87 35.90 26.30 33.33 14.84
5-shot 0.279 83.81 10.80 5.40 34.29 25.32 21.09 15.48
10-shot 0.256 82.35 11.15 6.50 30.77 23.38 16.67 15.81
after SFT zero-shot 0.217 84.62 11.72 3.66 25.00 22.40 15.65 12.26
Qwen3-14B-Base baseline zero-shot 0.292 81.52 15.22 3.26 23.72 16.56 39.12 19.35
1-shot 0.236 75.17 18.79 6.04 21.15 16.56 26.53 9.35
5-shot 0.190 73.22 23.01 3.77 18.59 14.61 13.27 10.65
10-shot 0.177 75.34 21.97 2.69 17.95 14.61 9.52 12.58
Qwen3-14B baseline zero-shot 0.340 79.67 17.52 2.80 32.37 22.08 45.92 11.94
1-shot 0.316 82.96 12.28 4.76 30.77 21.75 43.20 13.23
5-shot 0.263 85.20 12.99 1.81 31.09 25.32 26.19 9.68
10-shot 0.250 84.76 12.70 2.54 33.01 21.75 20.75 11.61
Qwen3-32B baseline zero-shot 0.270 81.23 16.72 2.05 25.00 24.35 26.53 14.84
1-shot 0.232 80.82 18.15 1.03 24.68 23.05 19.73 9.68
5-shot 0.201 81.10 18.50 0.39 24.36 19.48 14.97 8.39
10-shot 0.181 77.63 21.93 0.44 19.55 16.88 12.24 9.03
Table 25: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen3 family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using definition instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate